A triangle has two sides of length 11 and 19. The length of the third side must be between x and y, where x and y are as far apart as possible. What is the sum of x and y?
Answer by ikleyn(53996)

.

Due to the triangle inequalities, the third side should be longer than the difference (19-11 = 8)

and shorter than the sum (19 + 11 = 30).

So, in this problem x = 8, y = 30 (or, vice versa, x = 30, y = 11, the order does not matter).

In any case, the sum x+y is 8 + 30 = 38 = 30 + 8.

ANSWER. 38.